# Per-sender neural network classifiers for email authorship validation

### v1

@@ Abstract
Index Terms:
lateral spear phishing, business email compromise, email security, natural language processing, convolutional neural networks
1 Introduction
@@
The rest of the paper is organized as follows. Section 2 discusses prior research. Section 3 describes the construction of datasets for authorship validation. Section 4 details the Naive Bayes and Char-CNN classifiers used for authorship validation. Section 5 and 6 report and analyze the performance of the two classifiers on the constructed datasets. Section 7 describes how authorship validation classifiers may be incorporated into commercial email security systems. Section 8 summarizes our findings. Finally, 9 describes some ideas for continued exploration of authorship validation that are beyond the scope of this paper.
2 Related work
We review prior research work in three related fields: the use of sender profiles in detecting email-borne attacks, the use of neural networks and other machine learning (ML) techniques in identifying document authors, and publicly available descriptions of commercial email security systems. Our review reveals space for using per-sender neural network-based classification techniques in modern-day email security systems.
2.1 Sender profiles
@@
[10] explores sender-side detection of BEC emails (referred to as “spearphishing” in the cited paper). The paper’s authors describe a manual mechanism to build and maintain a behavioral profile per email sender in an organization. A profile captures a sender’s email habits with respect to a pre-determined list of attributes, including characters, words, and universal resource locators (URLs) in the body of the email, email formatting preferences (bullets, lists, paragraphs, signatures), email send times, and email recipients. When it receives a new email from a sender, the email security system extracts the above-mentioned features. Subsequently, the system feeds the features and the stored user profile to an ML, but a non-neural network classifier. In turn, the classifier provides a verdict regarding the new email’s authenticity compared to the sender’s stored profile.
@@
2.2 Authorship attribution
Several papers discuss the “authorship attribution” problem. Authorship attribution is the process of determining the author of a particular piece of writing. Authorship attribution can be formulated as either an open-class problem, where the true author might not be among a set of known authors, or a closed-class problem, where the true author is included in a finite set of authors.
[15] utilizes convolutional neural networks (CNNs) for authorship attribution, primarily focusing on learning patterns from character-level signals to identify an author’s style. The paper’s authors also investigated a novel architecture that combined character and word input channels within a CNN to capture stylistic and topical cues.
[16] presents a model for authorship attribution of short texts, specifically tweets, utilizing CNNs over character n-grams. The paper’s authors claim to be the first to present a CNN model based on character n-grams for this task and provide a comparison against traditional ML approaches.
[17] discusses advancements in using neural networks for character and word-level modeling for authorship attribution, particularly to address the scarcity of easily identifiable “signature” features in shorter texts.
[18] surveys several other papers on authorship attribution and related problems, including large language model (LLM) generated text detection and attribution.
While the authorship attribution literature is relevant, it solves a more difficult problem—identifying one author among several. In comparison, the authorship validation task, which is the subject of this paper, is conceptually more straightforward—it resembles a binary classification problem. In authorship validation, there is only one interesting author—we aim to judge whether this author actually penned the document (email). Simplifying the problem from authorship attribution to authorship validation makes the task computationally easier and cheaper and (potentially) enables commercial systems to employ neural networks per email sender.
2.3 Commercial email security systems
The following two papers provide insight into the design of email security systems.
@@ 2.3 Commercial email security systems
3 Dataset construction
@@ 3.2.1 LLM-generated inauthentic set
Appendix A.1 shows an example of a clean email from the dataset. Appendix A.2 shows the corresponding generated email. Most readers will notice the two emails’ differences in style and tone, even though the main point is identical.
@@
3.2.2 Other top senders’ emails as the inauthentic set
@@
Note that the emails in Inauthentic-2 are cleaned per the same tedious process described above in Section 3.1.1.
3.2.3 Mixture of emails as the inauthentic set
For the third inauthentic set (Inauthentic-3), we mixed the 369 LLM-generated emails in Section 3.2.1 (Inauthentic-1) with the 369 emails from the other top senders in Section 3.2.2 (Inauthentic-2), and an additional 279 BEC emails from a third-party [23], for a total of 1,017 inauthentic emails. Here, the idea is to simulate a heterogeneous combination of emails.
3.3 Creating datasets from authentic and inauthentic email sets
@@
4 Classifiers for authorship validation
We train two classifiers to assess the authenticity of email messages using the datasets described above in Section 3.
@@ 4.1 Naive Bayes
4.2 Char-CNN
We also employ a character-level convolutional neural network 00footnotemark: 0 to determine whether a specific individual authored a given document. The model performs binary classification, where a label of $1$ indicates authorship by the target individual and a label of 0 indicates otherwise. We train the network using binary cross-entropy loss.
00footnotetext: We refer to the character-level convolutional neural network implemented in this paper as Char-CNN.
@@ 4.2 Char-CNN
Each document is represented as a fixed-length sequence of character indices from a predefined vocabulary. The input tensor is defined as:
@@
where $B=32$ is the batch size and $L=1024$ is the maximum sequence length.
@@
Each character index is mapped to a dense vector using an embedding matrix:
@@
where $d=16$ is the embedding dimension. This tensor is permuted to:
$\mathbf{E}^{\prime}=\text{Permute}(\mathbf{E})\in\mathbb{R}^{B\times d\times L},$ (4)
to match the expected input shape of the convolutional layers.
@@
We apply three 1D convolutional layers with kernel sizes $k_{i}\in\{7,5,3\}$. For each kernel size $k_{i}$, we compute:
$\mathbf{C}_{i}=\text{BatchNorm}_{i}\left(\text{ReLU}\left(\text{Conv1D}_{i}(\mathbf{E}^{\prime})\right)\right)\in\mathbb{R}^{B\times F\times L_{i}},$ (5)
where $F=64$ is the number of filters and $L_{i}$ is the output sequence length for kernel size $k_{i}$.
Batch normalization is applied after each convolutional layer to improve training stability and accelerate convergence. The rectified linear unit (ReLU) activation function, defined as $\text{ReLU}(r)=\max(0,r)$, introduces non-linearity and helps mitigate the vanishing gradient problem.
@@
Each convolutional output is passed through adaptive max pooling over the sequence dimension:
@@
Adaptive max pooling reduces each feature map to a fixed output size by selecting the maximum value within dynamically sized regions. This allows the model to produce a consistent output shape regardless of input length.
Feature Concatenation and Dropout
The pooled features are concatenated and regularized using dropout:
$\mathbf{H}=\text{Dropout}([\mathbf{P}_{1};\mathbf{P}_{2};\mathbf{P}_{3}])\in\mathbb{R}^{B\times 3F}.$ (7)
Dropout is a regularization technique that randomly sets a fraction (here, $0.5$) of the input units to zero during training. This helps prevent overfitting by encouraging the network to learn redundant representations.
Fully Connected Layer and Output
The final layer is a fully connected layer followed by a sigmoid activation function:
$\mathbf{z}$ $=\mathbf{H}\cdot\mathbf{W}+\mathbf{b}\in\mathbb{R}^{B\times 1},$ (8)
$\hat{\mathbf{y}}$ $=\sigma(\mathbf{z})\in[0,1]^{B\times 1}.$ (9)
Loss Function
The model is trained to minimize the binary cross-entropy loss:
[…]
where $\mathbf{y}\in\{0,1\}^{B\times 1}$ is the true label vector and $\hat{\mathbf{y}}$ is the predicted probability of authentic authorship.
Predicted Label
During testing (evaluation), a document with a predicted probability of >=0.5>=0.5 is assigned a label of $1$. Otherwise, the document is assigned a label of 0.
5 Experimental results
Each of the four datasets described in Section 3 is split $80:20$ into training and test sets. Subsequently, both the Naive Bayes and Char-CNN classifiers are trained on the training set, and the classifiers’ performance is measured on the test set. Tables V, VI, VII, VIII show the performance results.
@@ 6 Analysis
6.1 Performance over Dataset-1
@@ 6.1.2 Char-CNN
Appendix A.3 shows an example of one of the false positive email files. While the email in the example is slightly more formal than the original, Char-CNN’s interpretation of the formality does not overcome the brevity of the message. Incidentally, Naive Bayes classified this email correctly.
Appendix A.4 shows an example of one of the false negative email files. This email contains a moderate-sized subject line and a single URL in the body. Given the modest dataset, it appears that Char-CNN did not have enough training examples to correctly adjudicate such emails. Naive Bayes also failed to classify this email correctly.
@@
6.2 Performance over Dataset-2
@@
6.3 Performance over Dataset-3
@@
6.4 Performance over Dataset-4
@@
7 The case for authorship validation in commercial systems
From Section 1, it is clear that a class of email-borne attacks—lateral spear phishing and lateral BEC—could be mitigated at the sender (potentially as the organization’s email handling system receives the email). From Section 2, we know that there is an opportunity to use ML-based classifiers in commercial email security systems to detect whether a sender actually authored an email that comes from their email account. From Sections 5 and 6, we know that even simple classifiers can provide an independent authorship validation signal and that there is room for neural network-based classifiers to improve the quality of that signal.
@@ 7.1 Sender profiles
7.2 Per-sender neural network classifiers
@@
7.3 Modular detection stack for sent emails
@@
7.4 Handling false positives and negatives
@@ 8 Conclusion
9 Future work
Several areas studied or briefly referenced in the paper invite future work:
  1. 1.
@@ 9 Future work
  3. 3.
    Implementation. Detailed design and real-life implementation of the sender detection system sketched out in Section 7.
  4. 4.
@@
Acknowledgment
We thank John Doe, an email security architect who has chosen to remain anonymous, for providing insights into the treatment of internal emails at large organizations. We also thank Swarup Selvaraman (email security expert), Minha Hwang, Paul Wells, and anonymous reviewers for discussion and feedback on earlier drafts of this paper
@@ Appendix A Email examples
A.1 Clean (authentic) email example
{spverbatim}
@@ A.1 Clean (authentic) email example
This is a resume of one guy I met in Houston a few months ago. He comes across as a very bright and qualified individual. He is interested in a position in London. Who is the best person in London to forward the resume to?
@@
A.2 Generated (inauthentic) email example
{spverbatim}

### v2

@@ Abstract
I Introduction
@@
The rest of the paper is organized as follows. Section II discusses prior research. Section III lays down the threat model assumed in this work. Section IV describes the construction of datasets for authorship validation. Section V details the Naive Bayes and Char-CNN classifiers used for authorship validation. Section VI and VII report and analyze the performance of the two classifiers on the constructed datasets. Section VIII describes how authorship validation classifiers may be incorporated into commercial email security systems. Section IX summarizes our findings. Finally, Section X describes some ideas for continued exploration of authorship validation that are beyond the scope of this paper.
II Related work
We review prior research work in three related fields: the use of sender profiles in detecting email-borne attacks, stylometry and authorship attribution, and publicly available descriptions of commercial email security systems. Our review reveals space for using per-sender neural network-based classification techniques in modern-day email security systems.
II-A Sender profiles
@@
[33] explores sender-side detection of BEC emails (referred to as “spearphishing” in the cited paper). The paper’s authors describe a manual mechanism to build and maintain a behavioral profile per email sender in an organization. A profile captures a sender’s email habits with respect to a pre-determined list of attributes, including characters, words, and universal resource locators (URLs) in the body of the email, email formatting preferences (bullets, lists, paragraphs, signatures), email send times, and email recipients. When it receives a new email from a sender, the email security system extracts the above-mentioned features. Subsequently, the system feeds the features and the stored user profile to an ML, but a non-neural network classifier. In turn, the classifier provides a verdict regarding the new email’s authenticity compared to the sender’s stored profile. Our paper can be thought of as a modern neural-network-based extension of the approach in [33].
@@
II-B Stylometry and authorship attribution
Stylometry studies measurable patterns of writing style to characterize authors or distinguish between them. Early work demonstrated that stable statistical properties of language can be used to resolve disputed authorship [22]. This analysis showed that function-word frequencies and other low-level stylistic cues persist across topics. These findings established a foundational statistical framework for later research in computational authorship analysis.
Distance-based stylometric methods were later formalized through the introduction of the Delta measure [3]. Delta quantifies stylistic distance using normalized word-frequency profiles and became widely adopted due to its simplicity. Subsequent work showed that character n-grams often outperform word-level features in capturing stylistic regularities. This result is particularly relevant for email, which is typically short, informal, and noisy.
A comprehensive survey of modern authorship attribution methods is provided in [32]. The survey organizes prior work by feature types, learning paradigms, and evaluation settings. It emphasizes that stylometric features are largely content-independent and complementary to semantic signals. These properties make stylometry well suited for security and forensic applications. Practical and computational aspects of authorship attribution are further examined in [15], including robustness across domains and adversarial settings.
Within this broader stylometric context, several papers focus specifically on the authorship attribution problem. Authorship attribution aims to determine the author of a given piece of writing and can be formulated as either an open-class or closed-class problem. Neural approaches have become increasingly prominent in this area, particularly those based on character-level modeling. Convolutional neural networks have been shown to effectively capture stylistic patterns from character-level signals [27]. Such models have also been extended to combine character and word representations to jointly model stylistic and topical cues.
Authorship attribution of short texts has received particular attention. Character n-gram CNNs have been successfully applied to short and noisy domains such as social media [31]. Subsequent work further explored neural architectures for character- and word-level modeling to address the lack of explicit stylistic “signature” features in short texts [18]. More recent surveys review a broad range of authorship-related problems, including attribution and large language model (LLM) generated text detection [14].
While closely related, authorship attribution addresses a more difficult problem than the one considered in this paper. Authorship attribution seeks to identify one author among many, whereas authorship validation focuses on verifying whether a single claimed author wrote a given document. Authorship validation can therefore be framed as a binary classification problem. This simplification reduces computational cost and makes it feasible to train sender-specific models for commercial email security systems.
Our work falls within the stylometric tradition by applying writing-style analysis to email security. We model per-sender writing style rather than message content, leveraging character-level representations that are widely recognized as effective stylometric features. By framing authorship validation as a behavioral signal for detecting impersonation and fraudulent messages, this work extends stylometric techniques to a practical email security setting.
II-C Commercial email security systems
Two publicly available papers provide insight into the design of email security systems.
@@ II-C Commercial email security systems
III Threat Model
This section defines the threat model assumed throughout this paper and makes explicit the security assumptions that motivate the authorship validation task. The structure of the threat model follows established threat modeling practices that emphasize assets, adversary capabilities, system boundaries, and defensive objectives [30, 29].
We consider an enterprise email environment in which users primarily trust internal email messages by default. We assume that standard email authentication mechanisms such as SPF, DKIM, and DMARC are already deployed and correctly enforce sender domain authenticity. As a result, the attacker cannot spoof external sender identities but can send emails from a legitimate internal account, a setting commonly assumed in prior email phishing threat models [23].
The primary adversary is an attacker who has obtained access to a valid internal email account through credential compromise, malware, or account takeover. The attacker’s goal is to impersonate the legitimate account owner in order to conduct lateral spear phishing or lateral business email compromise attacks. We assume the attacker seeks to exploit implicit trust relationships within the organization rather than relying on malware attachments or malicious links.
The attacker may send short, benign-looking emails that request sensitive actions, financial transfers, or confidential information. We assume the attacker has access to a limited subset of prior emails sent from the compromised account. This limitation may arise because large-scale mailbox exfiltration or prolonged access could trigger email security systems, data loss prevention mechanisms, or anomaly detection controls.
We further assume the attacker may use LLMs or other automated tools to generate emails that are grammatically correct and contextually appropriate. However, we assume the attacker can only condition such models on a relatively small number of authentic emails to avoid transferring large volumes of data or files. As a result, the attacker’s ability to closely mimic the legitimate sender’s writing style is imperfect and constrained.
The attacker’s capabilities include crafting messages that avoid common phishing indicators such as urgency cues, suspicious URLs, or obvious social engineering language. We assume the attacker does not have direct access to the parameters or internal state of the authorship validation classifier. We assume the attacker may adapt over time by observing which emails are blocked or flagged, but does not receive explicit classifier feedback beyond delivery success or failure.
From a machine learning security perspective, we assume a black-box evasion threat model in which the adversary can probe the system indirectly through email delivery outcomes but cannot inspect or manipulate model internals [25]. We assume that attacks targeting the confidentiality of model parameters or the integrity of the training pipeline are out of scope.
The primary asset protected by the system is the integrity of sender identity in internal email communication. A secondary asset is the trust relationship between employees that enables routine business operations, which has been identified as central to email-based attacks in prior threat modeling studies [23].
The system boundary includes the email subject and body content, which are the only inputs used by the authorship validation classifier. Email metadata, headers, attachments, and network-level signals are considered out of scope for the classifier studied in this paper.
We assume that the classifier operates as one signal within a larger, modular email security detection stack. The classifier’s output is not assumed to be a final enforcement decision, but rather an input into a broader aggregation framework, consistent with layered defense principles [30].
We explicitly scope our threat model to impersonation detection rather than intent classification. As a result, emails that are authentically authored but malicious in intent are considered out of scope. Similarly, emails that are inauthentic but benign in intent are still considered failures of sender identity integrity.
We assume the classifier is trained on historical emails authored by the legitimate sender and periodically retrained as new emails are sent. We assume the training data is predominantly clean and correctly labeled, though occasional mislabeled samples may occur.
We do not consider targeted data poisoning attacks against the training pipeline in this work. We do not consider denial-of-service attacks against the email infrastructure or classifier availability. The threat model assumes realistic operational constraints, including limited per-sender training data and modest computational budgets, in line with practical threat modeling guidance [29].
Under this model, the defender’s objective is to detect deviations from a sender’s established writing style that indicate possible impersonation. By making these assumptions explicit, the threat model clarifies the scope, limitations, and applicability of the proposed authorship validation system.
IV Dataset construction
@@ IV-B1 LLM-generated inauthentic set
Appendix A-A shows an example of a clean email from the dataset. Appendix A-B shows the corresponding generated email. Most readers will notice the two emails’ differences in style and tone, even though the main point is identical.
@@
IV-B2 Other top senders’ emails as the inauthentic set
@@
Note that the emails in Inauthentic-2 are cleaned per the same tedious process described above in Section IV-A1.
IV-B3 Mixture of emails as the inauthentic set
For the third inauthentic set (Inauthentic-3), we mixed the 369 LLM-generated emails in Section IV-B1 (Inauthentic-1) with the 369 emails from the other top senders in Section IV-B2 (Inauthentic-2), and an additional 279 BEC emails from a third-party [9], for a total of 1,017 inauthentic emails. Here, the idea is to simulate a heterogeneous combination of emails.
IV-C Creating datasets from authentic and inauthentic email sets
@@
V Classifiers for authorship validation
We train two classifiers to assess the authenticity of email messages using the datasets described above in Section IV.
@@ V-A Naive Bayes
V-B Character-Level Convolutional Neural Network (Char-CNN)
We employ a character-level convolutional neural network (Char-CNN) to determine whether a given document was authored by a specific individual. The task is formulated as binary classification, where a label of $1$ indicates authorship by the target individual and a label of 0 indicates non-authorship.
@@ V-B Character-Level Convolutional Neural Network (Char-CNN)
Each document is represented as a fixed-length sequence of character indices drawn from a predefined vocabulary. Documents longer than $L=1024$ characters are truncated, and shorter documents are padded with a special padding symbol. The resulting input tensor is defined as:
@@
where $B=32$ denotes the batch size.
@@
Character indices are mapped to dense vector representations using a learned embedding matrix:
@@
where $d=16$ is the embedding dimension. The embedding tensor is permuted to
$\mathbf{E}^{\prime}\in\mathbb{R}^{B\times d\times L},$ (4)
to match the input format expected by one-dimensional convolutional layers.
@@
We apply three parallel one-dimensional convolutional layers with kernel sizes $k\in\{7,5,3\}$. Each convolution produces $F=64$ feature maps. For each kernel size $k_{i}$, the convolutional transformation is defined as:
$\mathbf{C}_{i}=\text{ReLU}\!\left(\text{BatchNorm}_{i}\!\left(\text{Conv1D}_{i}(\mathbf{E}^{\prime})\right)\right),$ (5)
where batch normalization is applied to the convolutional outputs prior to the non-linear activation. This ordering stabilizes the distribution of intermediate activations and facilitates optimization.
@@
Each convolutional output is passed through adaptive max pooling over the temporal dimension:
@@
Adaptive max pooling aggregates each feature map into a single scalar by selecting the maximum activation across the sequence dimension.
Feature Concatenation and Regularization
The pooled feature vectors from all convolutional branches are concatenated and regularized using dropout:
$\mathbf{H}=\text{Dropout}\!\left([\mathbf{P}_{1};\mathbf{P}_{2};\mathbf{P}_{3}]\right)\in\mathbb{R}^{B\times 3F},$ (7)
where a dropout rate of $0.5$ is applied during training to reduce overfitting.
Output Layer
The concatenated feature representation is passed through a fully connected layer to produce a scalar logit:
$\mathbf{z}=\mathbf{H}\mathbf{W}+\mathbf{b}\in\mathbb{R}^{B\times 1},$ (8)
where $\mathbf{W}\in\mathbb{R}^{3F\times 1}$ and $\mathbf{b}\in\mathbb{R}$. No activation function is applied at this stage.
Training Objective
The model is trained using the binary cross-entropy loss with logits:
$\mathcal{L}(\mathbf{y},\mathbf{z})=\frac{1}{B}\sum_{b=1}^{B}\Big[-y_{b}\log\sigma(z_{b})-(1-y_{b})\log(1-\sigma(z_{b}))\Big],$ (9)
where $\sigma(\cdot)$ denotes the sigmoid function and $\mathbf{y}\in\{0,1\}^{B\times 1}$ represents the ground-truth labels. This formulation is implemented using BCEWithLogitsLoss, which improves numerical stability by combining the sigmoid operation with the loss computation.
Inference
During evaluation, predicted probabilities are obtained by applying the sigmoid function to the model logits. A document is classified as authored by the target individual if the predicted probability is greater than or equal to $0.5$; otherwise, it is classified as non-authored.
VI Experimental results
Each of the four datasets described in Section IV is split $80:20$ into training and test sets. Subsequently, both the Naive Bayes and Char-CNN classifiers are trained on the training set, and the classifiers’ performance is measured on the test set. Tables V, VI, VII, VIII show the performance results.
@@ VII Analysis
VII-A Performance over Dataset-1
@@ VII-A2 Char-CNN
Appendix A-C shows an example of one of the false positive email files. While the email in the example is slightly more formal than the original, Char-CNN’s interpretation of the formality does not overcome the brevity of the message. Incidentally, Naive Bayes classified this email correctly.
Appendix A-D shows an example of one of the false negative email files. This email contains a moderate-sized subject line and a single URL in the body. Given the modest dataset, it appears that Char-CNN did not have enough training examples to correctly adjudicate such emails. Naive Bayes also failed to classify this email correctly.
@@
VII-B Performance over Dataset-2
@@
VII-C Performance over Dataset-3
@@
VII-D Performance over Dataset-4
@@
VIII The case for authorship validation in commercial systems
From Section I, it is clear that a class of email-borne attacks—lateral spear phishing and lateral BEC—could be mitigated at the sender (potentially as the organization’s email handling system receives the email). From Section II, we know that there is an opportunity to use ML-based classifiers in commercial email security systems to detect whether a sender actually authored an email that comes from their email account. From Sections VI and VII, we know that even simple classifiers can provide an independent authorship validation signal and that there is room for neural network-based classifiers to improve the quality of that signal.
@@ VIII-A Sender profiles
VIII-B Per-sender neural network classifiers
@@
VIII-C Modular detection stack for sent emails
@@
VIII-D Handling false positives and negatives
@@ IX Conclusion
X Future work
Several areas studied or referenced in the paper invite future work. We aggregate these below for easy reference.
@@ Appendix A Email examples
A-A Clean (authentic) email example
@@
This is a resume of one guy I met in Houston a few months ago.
He comes across as a very bright and  qualified individual. He is interested in a position in London. Who is the best person in London to forward the resume to?
@@
A-B Generated (inauthentic) email example